OpenStudy (anonymous):

4xy + 8 = 36 A. (1,7) and (7,1) B. (1,7) and (4,9) C. (4,9) and (3,2) D. (7,1) and (3,2)

OpenStudy (anonymous):

It is A because 4*1*7=28 28+8=36 4*1*7+28 28+8=36

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Your welcome